OpenStudy (anonymous):

Find the value of x^3+2y^3 for x= 7 and y=1 Must show work.

OpenStudy (dumbcow):

this is known as plug and chug :)

OpenStudy (anonymous):

just plug in x and y so (7)^3 + 2(1)^3 7^3=343 2*1=2 2^3=8 so...343+8=351

OpenStudy (dumbcow):

@ses11 i dont think the "2" is attached to exponent

OpenStudy (anonymous):

@dumbcow oh yeah just noticed that :) so it would be 1^3 which is 1 and then 1*2=2 so 343+2=345